### v1

# Derivative-enhanced Deep Operator Network

Abstract
Deep operator networks (DeepONets), a class of neural operators that learn mappings between function spaces, have recently been developed as surrogate models for parametric partial differential equations (PDEs). In this work we propose a derivative-enhanced deep operator network (DE-DeepONet), which leverages the derivative information to enhance the prediction accuracy, and provide a more accurate approximation of the derivatives, especially when the training data are limited. DE-DeepONet incorporates dimension reduction of input into DeepONet and includes two types of derivative labels in the loss function for training, that is, the directional derivatives of the output function with respect to the input function and the gradient of the output function with respect to the physical domain variables. We test DE-DeepONet on three different equations with increasing complexity to demonstrate its effectiveness compared to the vanilla DeepONet.
Keywords: Neural operators, Operator learning, Derivative learning, DeepONet, dimension reduction, Adjoint method
1 Introduction
Using neural networks to approximate the maps between functions spaces governed by parametric PDEs can be very beneficial in solving many-query problems, typically arising from Bayesian inference, optimization under uncertainty, and Bayesian optimal experimental design. Indeed, once pre-trained on a dataset, neural networks are extremely fast to evaluate given unseen inputs, compared to traditional numerical methods like the finite element method. Recently many architectures are proposed to enhance the learning capacity, such as deep operator network (DeepONet) [11], physics-informed DeepONet [20], variational energy based DeepONet (V-DeepONet) [6], Fourier neural operator (FNO) [7], graph-based neural operator (GNO) [8], convolutional neural operator (CNO) [17]. Though these work demonstrate to be successful in approximating the output function, they do not necessarily provide accurate approximation of the derivative of the output with respect to the input, which typically needs to be computed many times in many problems in science and engineering such as PDE-constrained optimization problems for control, inference, and experimental design [13, 5].
In this paper, we propose a novel method to enhance the performance of DeepONet through derivative-based dimension reduction and the incorporation of derivative information in the training. For the dimension reduction, we examine two projection bases: Karhunen–Loève Expansion (KLE) basis and active subspace method (ASM) basis. We incorporate derivative information by introducing additional losses, enabling the model to learn the operator more effectively when training data are limited.
Our work is related to the derivative-informed projected neural network (DIPNet) [16] and derivative-informed neural operator (DINO) [15] and presents an extension to enhance the performance of the DeepONet. The notable differences include: Unlike the encoder-decoder architecture employed in their work, we use the DeepONet architecture in constructing the neural network. Additionally, while the output in their papers are considered as the observations of the solution at a finite number of points, our approach can approximate the full solution field. Another difference is that we also incorporate the spatial derivatives in the training.
Our main contributions can be summarized as follows: (1) We propose a novel method that significantly improves DeepONet’s approximation accuracy for the output function, its directional derivative with respect to the input function, as well as its spatial derivatives, especially when the training samples are limited. (2) We provide the details on the computation of derivative labels of the solution of PDEs in a general form as well as the derivative-based dimension reduction to largely reduce the computational cost. (3) We demonstrate the effectiveness (enhanced approximation accuracy and computation cost) of our method through experiments conducted on three distinct PDEs with increasing complexities.
The rest of the paper is presented as follows: In Section 2 we present preliminaries of the problem setup, high fidelity approximation, and DeepONet, followed by Section 3 on the detailed presentation of DE-DeepONet and demonstration experiments in Section 4. Discussion of the conclusion and limitation is presented in the last section.
2 Preliminaries
This section presents the problem setup, high-fidelity approximation using finite element, and the vanilla DeepONet.
2.1 Problem Setup
Let $\Omega\subset\mathbb{R}^{d}$ denote an open and bounded domain with boundary $\partial\Omega\subset\mathbb{R}^{d-1}$, where the dimension $d=1,2,3$. We consider a PDE of the general form defined in $\Omega$ as
$\mathcal{R}(m,u)=0,$ (1)
prescribed with proper boundary conditions. Here $m\in V^{\text{in}}$ is an input parameter function defined in a separable Banach space $V^{\text{in}}$ with probability measure $\nu$ and $u\in V^{\text{out}}$ is the output as the solution of the PDE defined in a separable Banach space $V^{\text{out}}$. Our goal is to construct a parametric model $\hat{u}(m;\theta)$ to approximate the solution operator that maps the parameter $m$ to the solution $u$.
Once constructed, the parametric model $\hat{u}(m;\theta)$ should be much more computationally efficient to evaluate compared to solving the PDE with high fidelity approximation.
2.2 High Fidelity Approximation
For the high fidelity approximation of the solution, we consider using a finite element method [14] in this work. We partition the domain $\Omega$ into a finite set of subregions, called cells or elements. Collectively, these cells form a mesh of the domain $\Omega$. Let $h$ represent the diameter of the largest cell. We denote $V^{\text{in}}_{h}$ indexed by $h$ as the finite element space for the approximation of the input space $V^{\text{in}}$ with Lagrange basis $\{\phi^{\text{in}}_{1},\cdots,\phi^{\text{in}}_{N^{\text{in}}_{h}}\}$ of dimension $N_{h}^{\text{in}}$ such that $\phi^{\text{in}}_{i}(x^{(j)})=\delta_{ij}$ at the finite element node $x^{(j)}$, with $\delta_{ij}$ being the Kronecker delta function. Similarly, we denote $V^{\text{out}}_{h}$ as the finite element space for the approximation of the solution space $V^{\text{out}}$ with basis $\{\phi^{\text{out}}_{1},\cdots,\phi^{\text{out}}_{N^{\text{out}}_{h}}\}$ of dimension $N_{h}^{\text{out}}$. Note that for the approximation to be high fidelity, $N_{h}^{\text{in}}$ and $N_{h}^{\text{out}}$ are often very large. To this end, we can write the high fidelity approximation of the input and output functions as
$m_{h}(x)=\sum_{i=1}^{N_{h}^{\text{in}}}m_{i}\phi_{i}^{\text{in}}(x)\;\text{ and }\;u_{h}(x)=\sum_{i=1}^{N_{h}^{\text{out}}}u_{i}\phi_{i}^{\text{out}}(x),$
with coefficient vectors $\bm{m}=(m_{1},\cdots,m_{N_{h}^{\text{in}}})^{T}\in\mathbb{R}^{N_{h}^{\text{in}}}$ and $\bm{u}=(u_{1},\cdots,u_{N_{h}^{\text{out}}})^{T}\in\mathbb{R}^{N_{h}^{\text{out}}}$, whose entries are the nodal values of $m$ and $u$ at the corresponding nodes.
2.3 Vanilla DeepONet
We briefly review the DeepONet architecture [11] with a focus on learning the solution operator of the PDE (1). To predict the evaluation of solution function $u$ at any point $x\in\Omega\cup\partial\Omega$ for any given input function $m$, [11] design a network architecture that comprises two separate neural networks: a trunk net $t(\cdot\;;\theta_{t})$, which takes the coordinate values of the point $x$ at which we want to evaluate the solution function as inputs, and a branch net $b(\cdot\;;\theta_{b})$, which takes the vector $\bm{m}$ encoding the parameter function $m$ as inputs. In [11], the vector $\bm{m}$ is the function evaluations at a finite set of fixed points $\{x^{(j)}\}_{j=1}^{N_{h}^{\text{in}}}\subseteq\Omega\cup\partial\Omega$, that is, $\bm{m}=(m(x^{(1)}),\cdots,m(x^{(N_{h}^{\text{in}})}))^{T}$, which corresponds to coefficient vector in the finite element approximation with Lagrange basis at the same nodes. If the solution function is scalar-valued, then both neural networks output vectors of the same dimensions. The prediction is obtained by taking the standard inner product between these two vectors and (optionally) adding a real-valued bias parameter, i.e,
$\hat{u}(\bm{m};\theta)(x)=\langle b(\bm{m};\theta_{b}),t(x;\theta_{t})\rangle+\theta_{\text{bias}},$
with $\theta=(\theta_{b},\theta_{t},\theta_{\text{bias}})$. If the solution $u$ is vector-valued of $N_{u}$ components, i.e., $u=(u^{(1)},\dots,u^{(N_{u})})$, as in our experiments, we can use one of the four approaches in [12] to construct the DeepONet. Specifically, for each solution component, we use the same outputs of branch net with dimension $N_{b}$ and different corresponding groups of outputs of trunk net to compute the inner product. More precisely, the solution $u^{(i)}$ of component $i=1,\dots,N_{u}$, is approximated by
$\hat{u}^{(i)}(\bm{m};\theta)(x)=\langle b(\bm{m};\theta_{b}),t^{(i)}(x;\theta_{t})\rangle+\theta_{\text{bias}}^{(i)},$ (2)
where $t^{(i)}(x;\theta_{t})=t(x;\theta_{t})[(i-1)N_{b}+1:iN_{b}]$, the vector slice of $t(x;\theta_{t})$ with indices ranging from $(i-1)N_{b}+1$ to $iN_{b}$, $i=1,\dots,N_{u}$.
For this construction, the outputs of the branch net can be interpreted as the coefficients of the basis learned through the trunk net. By partitioning the outputs of the trunk net into different groups, we essentially partition the basis functions used for predicting different components of the solution. The DeepONet is trained using dataset $\mathcal{D}=\{(m^{(i)},u^{(i)})\}_{i=1}^{N_{\mathcal{D}}}$ with $N_{\mathcal{D}}$ samples, where $m^{(i)}$ are random samples independently drawn from $\nu$ and $u^{(i)}$ are the solution of the PDE (with slight abuse of notation from the vector-valued solution) at $m^{(i)}$, $i=1,\dots,N_{\mathcal{D}}$.
3 DE-DeepONet
The vanilla DeepONet uses the input parameter and output solution pairs as the labels for the model training. The approximation for the derivative of the solution with repsect to the parameter (and the coordinate) are not necessarily accurate. However, in many downstream tasks such as Bayesian inference and experimental design, the derivatives are required. To better leverage our knowledge of the underlying PDE, we consider incorporating two additional derivative labels, that is, the Fréchet derivative $du(m;\cdot)$ and spatial derivative $\nabla_{x}u$, for the supervised training of the DeepONet, which we call derivative-enhanced DeepONet (DE-DeepONet). Let $\theta$ denote the trainable parameters in the DE-DeepONet. We propose to minimize the loss function
$\begin{split}L(\theta)&=\lambda_{1}\mathbb{E}_{m\sim\nu}\|u(m)-\hat{u}(m)\|_{L^{2}(\Omega)}^{2}\\ &+\lambda_{2}\mathbb{E}_{m\sim\nu}\|du(m;\cdot)-d\hat{u}(m;\cdot)\|_{\text{HS}}^{2}\\ &+\lambda_{3}\mathbb{E}_{m\sim\nu}\|\nabla_{x}u(m)-\nabla_{x}\hat{u}(m)\|_{L^{2}(\Omega)}^{2},\end{split}$ (3)
where the $L^{2}(\Omega)$ norm of a square integrable function $f$ is defined as $\|f\|_{L^{2}(\Omega)}^{2}=(\int_{\Omega}\|f(x)\|_{2}^{2}\;\mathrm{d}x)^{1/2}$, the Hilbert–Schmidt norm of an operator $T:H\to H$ that acts on a Hilbert space $H$ is defined as $\|T\|_{\text{HS}}^{2}=\sum_{i\in I}\|Te_{i}\|_{H}^{2}$, where $\{e_{i}\}_{i\in I}$ is an orthonormal basis of $H$. Here, $\lambda_{i}>0,i=1,2,3$ are hyperparameters that balance different loss terms.
The main challenge of minimizing the loss function (3) in practice is that with high fidelity approximation of the functions $m$ and $u$ using high dimensional vectors $\bm{m}$ and $\bm{u}$, the term $\|du(m;\cdot)-d\hat{u}(m;\cdot)\|_{\text{HS}}^{2}$ (approximately) becomes $\|\nabla_{\bm{m}}\bm{u}-\nabla_{\bm{m}}\bm{\hat{u}}\|_{F}^{2}$, where the Frobenius norm of a matrix $M\in\mathbb{R}^{m\times n}$ is defined as $\|M\|_{F}^{2}=\sum_{i=1}^{m}\sum_{j=1}^{n}M_{ij}^{2}$. It is a critical challenge to both compute and store the Jacobian matrix at each sample as well as to load and use it for training since it is often very large with size $N_{h}^{\text{out}}\times N_{h}^{\text{in}}$, with $N_{h}^{\text{in}},N_{h}^{\text{out}}\gg 1$.
To tackle this challenge, we employ dimension reduction for the high dimensional input vector $\bm{m}$. The reduced representation of $\bm{m}$ is given by projecting $\bm{m}$ into a low dimensional linear space spanned by a basis $\bm{\psi}_{1}^{\text{in}},\ldots,\bm{\psi}_{r}^{\text{in}}\in\mathbb{R}^{N_{h}^{\text{in}}}$, with $r\ll N_{h}^{\text{in}}$, that is,
$\widetilde{\bm{m}}=(\langle\bm{\psi}_{1}^{\text{in}},\bm{m}\rangle,\ldots,\langle\bm{\psi}_{r}^{\text{in}},\bm{m}\rangle)^{T}\in\mathbb{R}^{r}.$
To better leverage the information of both the input probability measure $\nu$ and the map $u:m\to u(m)$, we consider the basis generated by active subspace method (ASM) using derivatives [21]. ASM identifies directions in the input space that significantly affect the output variance, or in which the output is most sensitive. See Section 3.3 for the detailed construction. In this case, the term $\mathbb{E}_{m\sim\nu}\|du(m;\cdot)-d\hat{u}(m;\cdot)\|^{2}_{\text{HS}}$ can be approximated by $\frac{1}{N_{1}N_{2}}\sum_{i=1}^{N_{1}}\sum_{j=1}^{N_{2}}\|\nabla_{\widetilde{\bm{m}}}u(m^{(i)})(x^{(j)})-\nabla_{\widetilde{\bm{m}}}\hat{u}(m^{(i)})(x^{(j)})\|_{2}^{2}$ with a small amount of functions $m^{(i)}$ sampled from input probability measure $\nu$ and points $x^{(j)}$ in the domain $\Omega$. Note that $\nabla_{\widetilde{m}}\hat{u}(m^{(i)})(x^{(j)})$ is vector of size $r$, which is computationally feasible. For comparison, we also conduct experiments if the basis is the most commonly used Karhunen–Loève Expansion (KLE) basis.
We next formally introduces our DE-DeepONet, which uses dimension reduction for the input and incorporates two types of derivatives labels, that is, the directional derivative of output with respect to the input of the branch net and the derivative of output with respect to the input of the trunk net, as additional soft constraints into the loss function.
3.1 Model Architecture
We incorporate dimension reduction into DeepONet to construct a parametric model for approximating the solution operator. Specifically, if the solution is scalar-valued (the vector-valued case can be constructed similar to (2)), the prediction is given by
$\hat{u}(\bm{m};\theta)(x)=\langle b(\widetilde{\bm{m}};\theta_{b}),t(x;\theta_{t})\rangle+\theta_{\text{bias}},$ (4)
where $\theta=(\theta_{b},\theta_{t},\theta_{\text{bias}})$ are the parameters to be learned. The branch net $b(\cdot;\theta_{b})$ and trunk net $t(\cdot;\theta_{t})$ can be chosen as MLP, ResNet, etc. Note that the branch net takes a small vector of the projected parameter as input.
3.2 Loss Function
In practical training of the DE-DeepONet, we formulate the loss function as follows
$\begin{split}L(\theta)&=\frac{\lambda_{1}}{N_{\mathcal{D}}}\sum_{i=1}^{N_{\mathcal{D}}}\text{err}(\{(\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)}),u(m^{(i)})(x^{(j)}))\}_{j=1}^{N_{x}})\\ &+\frac{\lambda_{2}}{N_{\mathcal{D}}}\sum_{i=1}^{N_{\mathcal{D}}}\text{err}(\{(\nabla_{\bm{m}}\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)})\Psi^{\text{in}},\Phi^{\text{out}}(x^{(j)})(\nabla_{\bm{m}}\bm{u}(\bm{m}^{(i)})\Psi^{\text{in}})\}_{j=1}^{N_{x}})\\ &+\frac{\lambda_{3}}{N_{\mathcal{D}}}\sum_{i=1}^{N_{\mathcal{D}}}\text{err}(\{(\nabla_{x}\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)}),\nabla_{x}u(m^{(i)})(x^{(j)}))\}_{j=1}^{N_{x}}),\end{split}$ (5)
where $\{x^{(j)}\}_{j=1}^{N_{x}}$ are the nodes of the mesh, $\Psi^{\text{in}}=(\bm{\psi}_{1}^{\text{in}}|\cdots|\bm{\psi}_{r}^{\text{in}})$ is the matrix collecting the nodal values of the reduced basis of the input function space, and $\Phi^{\text{out}}(x)=(\phi^{\text{out}}_{1}(x),\ldots,\phi^{\text{out}}_{N_{h}^{\text{out}}}(x))$ is the vector-valued function consisting of the finite element basis functions of the output function space. The $\text{err}(\{(\bm{a}^{(i)},\bm{b}^{(i)})\}_{i=1}^{n})$ denotes the relative error between any two groups of vectors $\bm{a}^{(i)}$ and $\bm{b}^{(i)}$, computed as
$\text{err}(\{(\bm{a}^{(i)},\bm{b}^{(i)})\}_{i=1}^{n})=\frac{(\sum_{i=1}^{n}\|\bm{a}^{(i)}-\bm{b}^{(i)}\|_{2}^{2})^{1/2}}{\varepsilon+(\sum_{i=1}^{n}\|\bm{b}^{(i)}\|_{2}^{2})^{1/2}},$
where $\varepsilon>0$ is some small positive number to prevent the fraction dividing by zero.
In the following, we explain how to compute different loss terms in Equation 5
  • The first term is for matching the prediction of the parametric model $\hat{u}(\bm{m}^{(i)};\theta)$ evaluated at any set of points $\{x^{(j)}\}_{j=1}^{N_{x}}\subseteq\Omega\cup\partial\Omega$ with the high fidelity solution $u(m^{(i)})$ evaluated at the same points. The prediction $\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)})$ is straightforward to compute using Equation 4. This involves passing the reduced branch inputs $\widetilde{m}^{(i)}$ and the coordinates of point $x^{(j)}$ into the branch net and trunk net, respectively. The label $u(m^{(i)})(x^{(j)})$ is obtained using finite element method solvers.
  • The second term is for learning the directional derivative of the evaluation $u(x^{(j)})$ with respect to the input function $m$, in the direction of the reduced basis $\psi_{1}^{\text{in}},\ldots,\psi_{r}^{\text{in}}$. It can be shown that
    $\nabla_{\bm{m}}\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)})\Psi^{\text{in}}=\nabla_{\widetilde{\bm{m}}}\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)}).$
    Thus, the derivative of the outputs of the parametric model can be computed as the partial derivatives of the output with respect to the input of the branch net via automatic differentiation. On the other hand, the derivative labels
    $\Phi^{\text{out}}(x^{(j)})(\nabla_{\bm{m}}\bm{u}(m^{(i)})\Psi^{\text{in}})=(du(m^{(i)};\psi_{1}^{\text{in}})(x^{(j)}),\ldots,du(m^{(i)};\psi_{r}^{\text{in}})(x^{(j)}))$
    are obtained by first computing the Gateaux derivatives $du(m^{(i)};\psi_{1}^{\text{in}}),\ldots,du(m^{(i)};\psi_{r}^{\text{in}})$ and then evaluating them at $x^{(j)}$. See Section 3.4 for details about the computation of the Gateaux derivative $du(m;\psi)$.
  • The third term is for learning the spatial gradient of the output function $u$ at point $x^{(j)}$. The gradient outputs $\nabla_{x}\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)})$ are also computed via automatic differentiation, with the partial derivative being taken with respect to the input of the trunk net. The gradient labels $\nabla_{x}u(m^{(i)})(x^{(j)})$ are computed using the gradient of the finite element basis.
  • We initialize the loss weights $\lambda_{i}=1$ for $i=1,2,3$ and choose a loss balancing algorithm called the self-adaptive learning rate annealing algorithm [19] to update them at a certain frequency. This ensures that the gradient of each loss term have similar magnitudes, thereby enabling the neural network to learn all these labels simultaneously.
Remark.
[…]
3.3 Dimension Reduction
Throughout the paper, we assume that the parameter functions $m$ are independently drawn from some Gaussian random field [1]. In particular, we consider the case where the covariance function is the Whittle-Matern covariance function, that is, $m\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(\bar{m},\mathcal{C})$, where $\bar{m}$ is the (deterministic) mean function and $\mathcal{C}=(\delta I-\gamma\Delta)^{-2}$ ($I$ is the identity and $\Delta$ is the Laplacian) is an operator such that the square root of its inverse, $\mathcal{C}^{-\frac{1}{2}}$, maps random function $(m-\bar{m})$ to Gaussian white noise with unit variance [9]. The parameters $\delta,\gamma\in\mathbb{R}^{+}$ jointly control the marginal variance and correlation length.
We consider two well-known linear projection bases for dimension reduction of the parameter function.
Karhunen–Loève Expansion (KLE) basis.
The KLE basis is optimal in the sense that the mean-square error resulting from a finite representation of the random field $m$ is minimized [4]. It consists of eigenfunctions determined by the covariance function of the random field. Specifically, an eigenfunction $\psi$ of the covariance operator $\mathcal{C}=(\delta I-\gamma\Delta)^{-2}$ satisfies the differential equation
$\mathcal{C}\psi=\lambda\psi.$ (6)
When solved using the finite element method, Equation 6 is equivalent to the following linear system (See Section A.1 for the derivation.)
$M^{\text{in}}(A^{\text{in}})^{-1}M^{\text{in}}(A^{\text{in}})^{-1}M^{\text{in}}\bm{\psi}=\lambda M^{\text{in}}\bm{\psi},$ (7)
where the $(i,j)$-entries of matices $A^{\text{in}}$ and $M^{\text{in}}$ are given by
$A_{ij}^{\text{in}}=\delta\langle\phi_{j}^{\text{in}},\phi_{i}^{\text{in}}\rangle+\gamma\langle\nabla\phi_{j}^{\text{in}},\nabla\phi_{i}^{\text{in}}\rangle,\quad M_{ij}^{\text{in}}=\langle\phi_{j}^{\text{in}},\phi_{i}^{\text{in}}\rangle.$
[…]
Active Subspace Method (ASM) basis.
The active subspace method is a gradient-based dimension reduction method that looks for directions in the input space contributing most significantly to the output variability [3]. In contrast to the KLE basis, the ASM basis is more computationally expensive. However, since the ASM basis captures sensitivity information in the input-output map rather than solely the variability of the input space, it typically achieves higher accuracy in predicting the output than KLE basis. We consider the case where the output is a multidimensional vector [21], representing the nodal values of the output function $u$. The ASM basis $\psi$ satisfies the differential equation
$\mathcal{H}\psi=\lambda\mathcal{C}^{-1}\psi,$ (8)
where the action of operator $\mathcal{H}$ on function $\psi$ is given by
$\mathcal{H}\psi=\mathbb{E}_{m\sim\nu(m)}[d^{*}u(m;du(m;\psi))].$ (9)
Here, $du(m;\psi)$ is the Gateaux derivative of $u$ at $m\in V_{h}^{\text{in}}$ in the direction of $\psi\in V_{h}^{\text{in}}$, defined as $\lim_{\varepsilon\to 0}(u(m+\varepsilon\psi)-u(m))/\varepsilon$, and $d^{*}u(m;\cdot)$ is the adjoint of the operator $du(m;\cdot)$. When solved using finite element method,  Equation 8 is equivalent to the following linear system (See Section A.2 for the derivation.)
$H\bm{\psi}=\lambda C^{-1}\bm{\psi},$ (10)
where the action of matrix $H$ on vector $\bm{\psi}$ is given by
$H\bm{\psi}=\mathbb{E}_{\bm{m}\sim\nu(\bm{m})}[(\nabla_{\bm{m}}\bm{u})^{T}M^{\text{out}}(\nabla_{\bm{m}}\bm{u})\bm{\psi}],$ (11)
and the action of matrix $C^{-1}$ on vector $\bm{\psi}$ is given by
$C^{-1}\bm{\psi}=A^{\text{in}}(M^{\text{in}})^{-1}A^{\text{in}}\bm{\psi}.$ (12)
Here, $M^{\text{out}}$ denotes the mass matrix of the output function space, i.e., $M^{\text{out}}_{ij}=\langle\phi_{j}^{\text{out}},\phi_{i}^{\text{out}}\rangle$. In practice, when solving Equation 10, we obtain its left hand side through computing
$(\langle\mathcal{H}\psi,\phi_{1}^{\text{in}}\rangle,\ldots,\langle\mathcal{H}\psi,\phi_{N_{h}^{\text{in}}}^{\text{in}}\rangle)^{T}$
and its right hand side through the matrix-vector multiplication in Equation 12 (See  Section A.2 for details). Similar to the KLE case, let $\Psi^{\text{in}}$ denote the nodal values of $r$ dominant eigenfunctions. Since the eigenvectors $\bm{\psi}_{i}^{\text{in}}$ are $C^{-1}$-orthogonal, the reduced representation of $\bm{m}$ can be computed as $\widetilde{\bm{m}}=(\Psi^{\text{in}})^{T}C^{-1}\bm{m}$.
For the sake of scalability, we use the double pass randomized algorithm [18] in hIPPYlib (an open-source inverse problem python library) to solve the generalized eigenproblems Equation 7 and Equation 10.
3.4 Computation of the action of $\mathcal{H}$
By Equation 9, we use $N_{\text{grad}}$ samples $\{(m^{(i)},u^{(i)})\}_{i=1}^{N_{\text{grad}}}$ to compute the Monte Carlo estimate of the action of operator $\mathcal{H}$ on any function $\psi\in V_{h}^{\text{in}}$, that is,
$\mathcal{H}\psi\approx\frac{1}{N_{\text{grad}}}\sum_{i=1}^{N_{\text{grad}}}d^{*}u(m^{(i)};du(m^{(i)};\psi)).$ (13)
Each summation term in Equation 13 can be obtained by computing $p:=du(m;\psi)$ followed by $d^{*}u(m;p)$.
We first show how to compute $p$ given the PDE residual $\mathcal{R}(m,u)=0$. Since $u$ is uniquely determined by $m$, we can write $u=u(m)$ so that $\mathcal{R}(m,u(m))\equiv 0$ holds for any $m\in V_{h}^{\text{in}}$. Thus, for any $\varepsilon>0$ and $\psi\in V_{h}^{\text{in}}$, we have
$\mathcal{R}(m+\varepsilon\psi,u(m+\varepsilon\psi))=0.$
Using the Taylor expansion we obtain
$(\partial_{m}\mathcal{R}(m,u(m)))\varepsilon\psi+(\partial_{u}\mathcal{R}(m,u(m)))\delta u\approx 0,$ (14)
where $\delta u=u(m+\varepsilon\psi)-u(m)$. Dividing both sides of Equation 14 by $\varepsilon$ and letting $\varepsilon$ approach 00 yields
$(\partial_{m}\mathcal{R}(m,u(m)))\psi+(\partial_{u}\mathcal{R}(m,u(m)))du(m;\psi)=0.$
For ease of notation, we write
$\partial_{m}\mathcal{R}=\partial_{m}\mathcal{R}(m,u(m)),\quad\partial_{u}\mathcal{R}=\partial_{u}\mathcal{R}(m,u(m)).$
Then $p$ is the solution to the linear PDE
$(\partial_{m}\mathcal{R})\psi+(\partial_{u}\mathcal{R})p=0.$ (15)
We solve Equation 15 via its weak form
$\langle(\partial_{m}\mathcal{R})\psi,v\rangle+\langle(\partial_{u}\mathcal{R})p,v\rangle=0,$ (16)
where $v$ is a test function in $V_{h}^{\text{out}}$.
Next we show how to compute $d^{*}u(m;p)$. By Equation 15, we have $du(m;\psi)=-(\partial_{u}\mathcal{R})^{-1}(\partial_{m}\mathcal{R})\psi$. Thus,
$\begin{split}w:=d^{*}u(m;p)&=(-(\partial_{u}\mathcal{R})^{-1}(\partial_{m}\mathcal{R}))^{*}p\\ &=-(\partial_{m}\mathcal{R})^{*}(\partial_{u}\mathcal{R})^{-*}p.\end{split}$ (17)
Let $q:=(\partial_{u}\mathcal{R})^{-*}p$. We can solve for $p$ via the weak form
$\langle(\partial_{u}\mathcal{R})^{*}q,v\rangle=\langle p,v\rangle,$ (18)
where $v$ is a test function in $V_{h}^{\text{out}}$.
By Equation 17, we have $w=-(\partial_{m}\mathcal{R})^{*}q$. For any test function $v\in V_{h}^{\text{in}}$, it holds that
$\begin{split}\langle w,v\rangle&=\langle-(\partial_{m}\mathcal{R})^{*}q,v\rangle\\ &=\langle q,-(\partial_{m}\mathcal{R})v\rangle.\end{split}$ (19)
Note that we do not need to solve for $w$ explicitly; we only compute $\langle w,v\rangle$ with $v$ as the finite element basis functions $\phi_{1}^{\text{in}},\ldots,\phi_{N_{h}^{\text{in}}}^{\text{in}}$. The cost of computing the right hand side of Equation 19 arises from evaluating the directional derivative and its inner product with the finite element basis functions.
Remark.
When using the double pass randomized algorithm to obtain the first $r$ eigenpairs in  Equation 10, we need to compute the action of $\mathcal{H}$ on $2(r+s)$ random functions $V_{h}^{\text{in}}$ (e.g., their nodal values are sampled from the standard Gaussian distribution), where $s\in\mathbb{N}^{+}$ is typically a small oversampling parameter, often chosen between $5$ and $20$. To speed up the computation, we first compute the LU factorization of the matrices resulting from the discretization of the linear PDEs in Equation 16 and Equation 18. Then the action of $d^{*}u(m^{(i)};du(m^{(i)};\cdot))$ on these random functions can be efficiently computed via the forward and backward substitution. Furthermore, the computational time can be significantly reduced by parallelizing the computation of the average value in Equation 13 across multiple processors.
4 Experiments
In this section, we present experiment results to demonstrate the derivative-enhanced accuracy and cost of our method on three test PDEs, including both linear and nonlinear, both scalar-valued and vector-valued solutions. Details about the data generation and training can be found in Appendix B.
4.1 Input Probability Measure
In all test cases, we assume that the input functions $m^{(i)}$ are i.i.d. samples drawn from a Gaussian random field with mean function $\bar{m}$ and covariance operator $(\delta I-\gamma\Delta)^{-\alpha}$, with $\alpha=2$ so the covariance operator is of trace class. It is worth noting that the parameters $\gamma$ and $\delta$ jointly control the marginal variance and correlation length, for which we take values to have large variation of the samples so that it is difficult for the neural network approximation with relatively large errors. We use a scalable sampling algorithm in hIPPYlib [18] to generate the samples.
4.2 Governing Equations
We consider a linear and scalar-valued Darcy flow equation, a nonlinear vector-valued hyperelasticity equation with one state (displacement), and a nonlinear vector-valued Navier–Stokes equations with multiple states (velocity and pressure). Specifically, we learn the map from (the logarithm of) the diffusion coefficient field to the solution of the Darcy flow equation with zero Dirichlet boundary condition. For the hyperelasticity equation, we consider an experimental scenario where a square of hyperelasticity material is secured along its left edge while a fixed upward-right force is applied to its right edge [2]. Our goal is to learn the map between (the logarithm of) the material’s Young’s modulus and its displacement. Lastly, we consider the Navier–Stokes equations that describes viscous, incompressible fluid flow. In particular, we consider the lid-driven cavity case where a square cavity consisting of three rigid walls with no-slip conditions and a lid moving with a tangential unit velocity. We consider that the uncertainty comes from the viscosity term and aim to predict the velocity field.
Darcy flow equation.
The strong form of the PDE can be written as
$\begin{cases}-\nabla\cdot(e^{m}\nabla u)=1,\quad x\in\Omega,\\ u(x)=0,\quad x\in\partial\Omega.\end{cases}$
Here, we assume that the randomness arises from the viscosity term $e^{m}$. For demonstration, we choose $\bar{m}=0$, $\delta=1.0$ and $\gamma=0.1$. Note that the PDE is linear but the solution operator is not. See Figure 1 for the visualization for input-output pairs of the Darcy flow equation.
Figure 1: Visualization of input-output pairs of Darcy flow equation.
Hyperelasticity equation.
We follow the problem setup in [2]. Write $X$ (instead of $x$) for a material point in the domain $\Omega$ and $u=u(X):\mathbb{R}^{2}\to\mathbb{R}^{2}$ for the displacement of the material point. Under the influence of internal and/or external forces, the material point is mapped to a spatial point $x=x(X)=X+u(X):\mathbb{R}^{2}\to\mathbb{R}^{2}$. Let $F=\nabla_{X}x=I+\nabla_{X}u:\mathbb{R}^{2}\to\mathbb{R}^{2\times 2}$ denote the deformation gradient. For a hyperelasticity material, the internal forces can be derived from a strain energy density
$W(X,C)$ $=\frac{\mu(X)}{2}(\mathrm{tr}(C)-3)+\frac{\lambda(X)}{2}(\ln(J))^{2}-\mu(X)\ln(J).$
Here, $C=F^{T}F$ is the right Cauchy-Green strain tensor, $\mathrm{tr}(C)$ is the trace of matrix $C$, $J$ is the determinant of matrix $F$, and $\mu(X),\lambda(X):\mathbb{R}^{2}\to\mathbb{R}$ are the Lamé parameters which we assume to be related to Young’s modulus of elasticity $E(X):\mathbb{R}^{2}\to\mathbb{R}$ and Poisson ratio $\nu\in\mathbb{R}$
$\mu(X)=\frac{E(X)}{2(1+\nu)},\quad\lambda(X)=\frac{\nu E(X)}{(1+\nu)(1-2\nu)}.$
We assume the randomness comes from the Young’s modulus $E(X)=e^{m(X)}+1$. Let $S=2\frac{\partial W}{\partial C}$ denote the second Piola-Kirchhoff stress tensor. We consider the case where the left boundary of the material is fixed, and the right boundary is subjected to stretching by an external force $t=t(X):\mathbb{R}^{2}\to\mathbb{R}^{2}$. The strong form of the steady state PDE can be written as
$\begin{cases}\nabla_{X}\cdot(FS)=0,&\quad X\in\Omega,\\ u=0,&\quad X\in\Gamma_{\text{left}},\\ FS\cdot n=0,&\quad X\in\Gamma_{\text{top}}\cup\Gamma_{\text{bottom}},\\ FS\cdot n=t,&\quad X\in\Gamma_{\text{right}},\end{cases}$
where $\Gamma_{\text{left}},\Gamma_{\text{right}},\Gamma_{\text{top}}$ and $\Gamma_{\text{bottom}}$ denote the left, right, top, and bottom boundary of the material domain $\Omega$, respectively, and $n$ is the unit outward normal vector on the boundary. Our goal is to learn the operator that maps the parameter $m$ to the displacement $u$. For demonstration, we choose $\bar{m}=0$, $\delta=0.4$, $\gamma=0.04$, Poisson ratio $\nu=0.4$, and the external force
$t(X)=\left(0.06\exp(-0.25|X_{2}-0.5|^{2}),0.03(1+0.1X_{2})\right)^{T}.$
In practice, we solve the PDE by first formulating the energy $\widetilde{W}$ in the weak form
$\widetilde{W}=\int_{\Omega}W\mathrm{d}X-\int_{\Gamma_{\text{right}}}\langle t,u\rangle\;\mathrm{d}s$
and then solving for $u$ that satisfies the stationary condition, that is, the equation
$d\widetilde{W}(u;v)=0,$
holds for any test function $v$ in the state space. See Figure 2 for the visualization for input-output pairs of the hyperelasticity equation.
Figure 2: Visualization of input-output pairs of hyperelasticity equation.
Navier–Stokes equation.
Let $u=u(x)\in\mathbb{R}^{2}$ and $p=p(x)\in\mathbb{R}$ denote the velocity and pressure at point $x\in\Omega=(0,1)^{2}$. The strong form of the PDE can be written as
$\begin{cases}-\nabla\cdot e^{m}\nabla u+(u\cdot\nabla)u+\nabla p=0,&\quad x\in\Omega,\\ \nabla\cdot u=0,&\quad x\in\Omega,\\ u=(1,0)^{T},&\quad x\in\Gamma_{\text{top}},\\ u=(0,0)^{T},&\quad x\in\Gamma\setminus\Gamma_{\text{top}},\end{cases}$
where $\Gamma_{\text{top}}$ and $\Gamma$ denote the left and whole boundary of the cavity domain $\Omega$, respectively. Here, we assume that the randomness arises from the viscosity term $e^{m}$. Our goal is learn the operator that maps the parameter $m$ to the velocity $u$. For demonstration, we choose $\bar{m}=6.908$ ($e^{\bar{m}}\approx 10^{3}$), $\delta=0.6$, and $\gamma=0.06$. See Figure 3 for the visualization for input-output pairs of the Navier–Stokes equations.
Figure 3: Visualization of input-output pairs of Navier–Stokes equations.
4.3 Evaluation Metrics
We use the following three metrics to evaluate the performance of different methods. For $u(m)(x)$ approximations, we compute the relative error in the $L^{2}(\Omega)$ norm and $H^{1}(\Omega)$ norm on the test dataset $\mathcal{D}_{\text{test}}=\{(m^{(i)},u^{(i)})\}_{i=1}^{N_{\text{test}}}$, that is,
$\frac{1}{N_{\text{test}}}\sum_{m^{(i)}\in\mathcal{D}_{\text{test}}}\frac{\|\hat{u}(m^{(i)};\theta)-u(m^{(i)})\|_{L^{2}(\Omega)}}{\|u(m^{(i)})\|_{L^{2}(\Omega)}},\quad\frac{1}{N_{\text{test}}}\sum_{m^{(i)}\in\mathcal{D}_{\text{test}}}\frac{\|\hat{u}(m^{(i)};\theta)-u(m^{(i)})\|_{H^{1}(\Omega)}}{\|u(m^{(i)})\|_{H^{1}(\Omega)}},$
where
$\|u\|_{L^{2}(\Omega)}=\int_{\Omega}\|u(x)\|^{2}_{2}\;\mathrm{d}x=\bm{u}^{T}M^{\text{out}}\bm{u},\quad\|u\|_{H^{1}(\Omega)}=(\|u\|_{L^{2}(\Omega)}^{2}+\|\nabla u\|_{L^{2}(\Omega)}^{2})^{1/2}.$
For $du(m;\cdot)(x)$ approximations, we compute the relative error in the Frobenius norm on $\mathcal{D}_{\text{test}}$ along random directions $\omega=\{\omega_{i}\}_{i=1}^{N_{\text{dir}}}$, that is,
$\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{m^{(i)}\in\mathcal{D}_{\text{test}}}\frac{\|d\hat{u}(m^{(i)};\omega)(x)-du(m^{(i)};\omega)(x)\|_{F}}{\|du(m^{(i)};\omega)(x)\|_{F}},$
where $\omega_{i}$ are samples drawn from the same distribution as $m$.
4.4 Main Results
We observe consistent performance in terms of enhanced approximation accuracy for the three different equations. Here we only present a part of the results for the Navier–Stokes equations and leave all other results for other equations, including also hyperparameter tuning in the training of the neural networks, the architecture of different neural networks, the configuration of data generations, the full set of test errors, visualization of the KLE and ASM basis functions, and visualization of the prediction of both the solution and derivatives in Appendix B. We show the tradeoff in terms of prediction error and computational cost of different methods. We find that the advantages of our method are especially evident for the more complex and challenging Navier–Stokes equations.
Test errors.
In Figure 4, we show a comparison of different methods for predicting the velocity field in the $y$ direction and its derivative with respect to parameter in the direction of $\{\omega_{i}\}_{i=1}^{N_{\text{dir}}}$ with $N_{\text{dir}}=128$. We observe that the reduced basis DeepONet, even without using derivative labels for training, has a prediction error comparable to that of the vanilla DeepONet. This suggests that the chosen number of reduced bases $(r=16)$ is sufficiently large to predict a high-quality output function.
We also find that incorporating two additional derivative losses (denoted as ‘dm’ loss and ‘dx’ loss) into the training of the reduced basis DeepONet significantly lower the errors (to between one-thirds and one-half of their original values in predicting the velocity field in the $y$ direction) in predicting both the solution $u$ and its directional derivative with respect to the input function $m$. See Figure 5 and Figure 6 for a visualization comparing the prediction of velocity-y and its directional derivative, where the DE-DeepONet with ASM+dm+dx using the two derivative labels provides a more accurate prediction.
In addition, we conduct experiments for comparing the results of adding solely the ‘dm’ loss versus adding both ‘dm’ and ‘dx’ losses. The details are provided in Section B.3. We observe that incorporating ‘dm’ loss makes a dominant contribution to error reduction, while incorporating ‘dx’ loss has a relatively smaller effect on reducing the errors.
Refer to caption Refer to caption Refer to caption
Figure 4: Mean relative errors over 5 trials with a varying number of training samples for the Navier–Stokes equations using different methods. Relative errors in the $L^{2}(\Omega)$ norm (left) and $H^{1}({\Omega})$ norm (middle) for the prediction of the velocity-y. Right: Relative error in the Frobenius (Fro) norm for the prediction of the derivative of the velocity-y with respect to the parameter in the direction of the 128 samples drawn from the same distribution as parameter.
Figure 5: Comparison of the predictions of velocity-y with (16, 64, 256, 1024) training samples using different methods. Top: Vanilla DeepONet; Middle: ASM; Bottom: DE-DeepONet using ASM + dm + dx.
Figure 6: Comparison of the predictions of directional derivative $du_{2}(m;\omega_{1})$ with (16, 64, 256, 1024) training samples using different methods. Top: Vanilla DeepONet; Middle: ASM; Bottom: DE-DeepONet using ASM + dm + dx.
Computational cost.
We use MPI and the finite element library FEniCS [10] to distribute the computational load of offline data generation to 64 CPUs for all the PDEs considered in this work. See Table 1 for the wall clock time for generating the samples as Gaussian random fields (GRFs), solving the PDEs, computing the $r=16$ reduced basis functions (KLE or ASM) corresponding to the $16$ largest eigenvalues, generating the derivative labels dm and dx with respect to the parameter and space, respectively. Note that generating the derivative labels is more expensive than generating the PDE solutions, but not by large, especially for the more complex and expensive Navier–Stokes equations.
Table 1: Wall clock time (in seconds) for the data generation on 2 X AMD EPYC 7543 32-Core Processors
Dataset Process
GRFs PDEs KLE ASM dm labels dx labels
$(N_{\text{all}}=2000)$ $(N_{\text{all}}=2000)$ $(r=16)$ $(r=16)$ $(N_{\text{all}}=2000,r=16)$ $(N_{\text{all}}=2000)$
($64$ CPUs) ($64$ CPUs) ($1$ CPU) ($16$ CPUs) ($64$ CPUs) ($64$ CPUs)
Darcy flow $1.1$ $1.4$ $0.4$ $1.1$ $8.3$ $4.6$
Hyperelasticity $1.1$ $9.7$ $0.4$ $1.4$ $19.5$ $8.8$
Navier–Stokes $1.9$ $99.1$ $1.3$ $9.7$ $125.5$ $45.8$
We present comparisons of the computational cost of model training using different methods with two different number of training samples ($N_{\text{train}}=64$, corresponding to a limited training dataset setting, and $N_{\text{train}}=1024$, corresponding to a sufficient training dateset setting) in Table 2. We find that incorporating derivative losses leads to larger computational cost as expected. However, when the training data are limited, the increased computation cost is compensated for a significant reduction of errors. We note that there are potential ways to reduce the cost, e.g., by training the model with additional derivative losses only during the later stage of training, or by using fewer points for computing the derivative losses in each iteration. Additionally, the number of trainable parameters in the neural network is shown in Table 3. We observe that the reduced basis DeepONet has an order of magnitude fewer parameters than the vanilla DeepONet. Thanks to the dimension reduction in the input parameter, we are able to efficiently compute and store the derivatives as well as train the reduced basis DeepONet.
Table 2: Wall clock time (in seconds per epoch) for the neural network training on a single NVIDIA RTX A6000 GPU
Dataset Method ($N_{\text{train}}=64$, $N_{\text{train}}=1024$)
Vanilla Reduced Reduced + dm Reduced + dm + dx
Darcy flow $(0.04,0.64)$ $(0.03,0.50)$ $(0.09,1.58)$ $(0.14,2.39)$
Hyperelasticity $(0.04,0.73)$ $(0.03,0.51)$ $(0.14,2.32)$ $(0.18,3.02)$
Navier–Stokes $(0.04,0.75)$ $(0.04,0.64)$ $(0.15,2.59)$ $(0.22,3.60)$
Table 3: Number of trainable parameters in the neural network
Dataset Neural network size
Vanilla Reduced
Darcy flow $3,420,417$ $101,633$
Hyperelasticity $3,716,482$ $249,858$
Navier–Stokes $3,716,482$ $496,770$
5 Discussion
In this work, we proposed a new neural operator–Derivative-enhanced Deep Operator Network (DE-DeepONet). Compared to vanilla DeepONet, DE-DeepONet receives a reduced representation of input function and has two additional loss terms, which are for the supervised learning of the derivative of the output with respect to the inputs of both the branch net and the trunk net. Our experiments demonstrate that adding these two loss terms to the loss function greatly enhances the accuracy of both function approximations and derivative approximations, especially when the training data are limited.
We observe that when the training data are large enough, the benefits brought by adding derivative loss terms become less significant (that is, with more and more training data, we expect a relatively smaller increase in marginal accuracy and a relatively larger increase in marginal computational cost). However, in computational science and engineering, unlike other machine learning fields such as computer vision, training data in complex and nonlinear physical system applications are often scarce due to the high costs of computational simulations. We demonstrated that the advantage of our method becomes more apparent for the more complex Navier–Stokes equations which require more iterations in the nonlinear solves than the hyperelasticity equation.
Appendix A Proofs
A.1 Proof of the Equivalence of Equation 6 and Equation 7
By the definition of $\mathcal{C}=(\delta I-\gamma\Delta)^{-2}$, Equation 6 is equivalent to
$\psi=\lambda(\delta I-\gamma\Delta)^{2}\psi.$ (20)
We first compute $(\delta I-\gamma\Delta)\psi$. To do this, let $p=(\delta I-\gamma\Delta)\psi$ and multiply both sides of this equation by a test function $v$ and integrate
$\int_{\Omega}p(x)v(x)\;\mathrm{d}x$ $=\int_{\Omega}(\delta I-\gamma\Delta)\psi(x)v(x)\;\mathrm{d}x$
$=\delta\int_{\Omega}\psi(x)v(x)\;\mathrm{d}x+\gamma\int_{\Omega}\langle\nabla\psi(x),\nabla v(x)\rangle\;\mathrm{d}x,$ (21)
where, in the second equality, we use integration by parts and the assumption that the test function vanishes on the boundary. Then we substitute $v$ with all of the finite element basis functions $\phi_{1}^{\text{in}},\ldots,\phi_{N_{h}^{\text{in}}}^{\text{in}}$ and collect the corresponding linear equations for the nodal values of $p$
$A^{\text{in}}\bm{\psi}=M^{\text{in}}\bm{p},$ (22)
where the $(i,j)$-entries of $A^{\text{in}}$ and $M^{\text{in}}$ are given by
$A_{ij}^{\text{in}}=\delta\langle\phi_{j}^{\text{in}},\phi_{i}^{\text{in}}\rangle+\gamma\langle\nabla\phi_{j}^{\text{in}},\nabla\phi_{i}^{\text{in}}\rangle,\quad M_{ij}^{\text{in}}=\langle\phi_{j}^{\text{in}},\phi_{i}^{\text{in}}\rangle.$
By Equation 20, function $p$ satisfies $\psi=\lambda(\delta I-\gamma\Delta)p$. In the same manner we can see that
$\lambda A^{\text{in}}\bm{p}=M^{\text{in}}\bm{\psi}.$ (23)
Note that both matrices $M^{\text{in}}$ and $A^{\text{in}}$ are symmetric positive definite and thus nonsingular. Combining Equation 22 with Equation 23 yields
$\lambda A^{\text{in}}(M^{\text{in}})^{-1}A^{\text{in}}\bm{\psi}=M^{\text{in}}\bm{\psi},$ (24)
or equivalently,
$M^{\text{in}}(A^{\text{in}})^{-1}M^{\text{in}}(A^{\text{in}})^{-1}M^{\text{in}}\bm{\psi}=\lambda M^{\text{in}}\bm{\psi}.$ (25)
A.2 Proof of the Equivalence of Equation 8 and Equation 10
By Equation 8, for any test function $v\in V_{h}^{\text{in}}$, it holds that
$\langle\mathcal{H}\psi,v\rangle=\langle\lambda\mathcal{C}^{-1}\psi,v\rangle.$ (26)
In particular, for any $m\sim\nu(m)$, as we let $v$ go through all of the finite element basis functions $\phi_{i}^{\text{in}}\in V_{h}^{\text{in}}$, we can show that
$\begin{pmatrix}\langle d^{*}u(m;du(m;\psi)),\phi_{1}^{\text{in}}\rangle\\ \vdots\\ \langle d^{*}u(m;du(m;\psi)),\phi_{N_{h}^{\text{in}}}^{\text{in}}\rangle\\ \end{pmatrix}=(\nabla_{\bm{m}}\bm{u})^{T}M^{\text{out}}(\nabla_{\bm{m}}\bm{u})\bm{\psi}.$ (27)
Indeed, by the definition of Gateaux derivative, we have
$\begin{split}du(m;\psi)&=\lim_{\varepsilon\to 0}\frac{u(m+\varepsilon\psi)-u(m)}{\varepsilon}\\ &=\lim_{\varepsilon\to 0}\sum_{i=1}^{N_{h}^{\text{out}}}\frac{\bm{u}_{i}(m+\varepsilon\psi)-\bm{u}_{i}(m)}{\varepsilon}\phi_{i}^{\text{out}}(x)\\ &=\lim_{\varepsilon\to 0}\sum_{i=1}^{N_{h}^{\text{out}}}\frac{\bm{u}_{i}(\bm{m}_{1}+\varepsilon\bm{\psi}_{1},\cdots,\bm{m}_{N_{h}^{\text{in}}}+\varepsilon\bm{\psi}_{N_{h}^{\text{in}}})-\bm{u}_{i}(\bm{m}_{1},\cdots,\bm{m}_{N_{h}^{\text{in}}})}{\varepsilon}\phi_{i}^{\text{out}}(x)\\ &=\sum_{i=1}^{N_{h}^{\text{out}}}(\frac{\partial\bm{u}_{i}}{\partial\bm{m}_{1}}\bm{\psi}_{1}+\cdots+\frac{\partial\bm{u}_{i}}{\partial\bm{m}_{N_{h}^{\text{in}}}}\bm{\psi}_{N_{h}^{\text{in}}})\phi_{i}^{\text{out}}(x)\\ &=\phi^{\text{out}}(x)(\nabla_{\bm{m}}\bm{u})\bm{\psi},\end{split}$ (28)
where $\phi^{\text{out}}(x)=(\phi_{1}^{\text{out}}(x),\ldots,\phi_{N_{h}^{\text{out}}}(x))$ are the finite element basis functions of output function space $V_{h}^{\text{out}}$.
Then for any test function $v\in V_{h}^{\text{in}}$, it holds that
[…]
$=\bm{v}^{T}(\nabla_{\bm{m}}\bm{u})^{T}M^{\text{out}}(\nabla_{\bm{m}}\bm{u})\bm{\psi},$
where $M^{\text{out}}$ is the mass matrix of output function space $V_{h}^{\text{out}}$, i.e., $M^{\text{out}}_{ij}=\langle\phi_{j}^{\text{out}},\phi_{i}^{\text{out}}\rangle$ for $1\leq i,j\leq N_{h}^{\text{out}}$. Note that if we replace $v$ by the $i$-th finite element basis functions $\phi_{i}^{\text{in}}$, then $\bm{v}$ becomes the standard unit vector $e_{i}\in\mathbb{R}^{N_{h}^{\text{in}}}$ (with the $k$-th entry one and all others zero). Thus,
$\langle d^{*}u(m;p),\phi_{i}^{\text{in}}\rangle=e_{i}^{T}(\nabla_{\bm{m}}\bm{u})^{T}M^{\text{out}}(\nabla_{\bm{m}}\bm{u})\bm{\psi},\quad 1\leq i\leq N_{h}^{\text{in}}.$
Concatenating all the above equations yields
$\begin{pmatrix}\langle d^{*}u(m;p),\phi_{1}^{\text{in}}\rangle\\ \vdots\\ \langle d^{*}u(m;p),\phi_{N_{h}^{\text{in}}}^{\text{in}}\rangle\\ \end{pmatrix}=(\nabla_{\bm{m}}\bm{u})^{T}M^{\text{out}}(\nabla_{\bm{m}}\bm{u})\bm{\psi}.$
Next, we prove that
$\begin{pmatrix}\langle\lambda\mathcal{C}^{-1}\psi,\phi_{1}^{\text{in}}\rangle\\ \vdots\\ \langle\lambda\mathcal{C}^{-1}\psi,\phi_{N_{h}^{\text{in}}}^{\text{in}}\rangle\\ \end{pmatrix}=\lambda A^{\text{in}}(M^{\text{in}})^{-1}A^{\text{in}}\bm{\psi}.$ (29)
Indeed, if we let $w=\lambda\mathcal{C}^{-1}\psi$, then similar to the argument in Section A.1, we have
$\lambda A^{\text{in}}(M^{\text{in}})^{-1}A^{\text{in}}\bm{\psi}=M^{\text{in}}\bm{w},$
Note that
$\langle w,\phi_{i}^{\text{in}}\rangle=e_{i}^{T}M^{\text{in}}\bm{w},\quad 1\leq i\leq N_{h}^{\text{in}}.$
Thus,
$\begin{pmatrix}\langle w,\phi_{1}^{\text{in}}\rangle\\ \vdots\\ \langle w,\phi_{N_{h}^{\text{in}}}^{\text{in}}\rangle\\ \end{pmatrix}=IM^{\text{in}}\bm{w}=\lambda A^{\text{in}}(M^{\text{in}})^{-1}A^{\text{in}}\bm{\psi}.$
Combining (26),  (27) and (29) yields
$\mathbb{E}_{\bm{m}\sim\nu(\bm{m})}[(\nabla_{\bm{m}}\bm{u})^{T}M^{\text{out}}(\nabla_{\bm{m}}\bm{u})\bm{\psi}]=\lambda A^{\text{in}}(M^{\text{in}})^{-1}A^{\text{in}}\bm{\psi}.$
Appendix B Experimental Details
B.1 Data generation
For all PDEs in this work, we use the class dolfin.UnitSquareMesh to create a triangular mesh of the 2D unit square with $64$ cells in horizontal direction and $64$ cells in vertical direction. For the Darcy flow equation and hyperelasticity equation, we set the direction of the diagonals as ’right’, while for the Navier–Stokes equation, we set the direction of the diagonals as ’crossed’. See Figure 7 for a visualization of the unit square mesh with a $10$ by $10$ resolution.
Figure 7: Visualization of the $10$ by $10$ unit square mesh. Left: diagonal=‘right’; Right: diagonal=‘crossed’
We use the class dolfin.FunctionSpace or dolfin.VectorFunctionSpace to create the finite element function space of input function and output function. For the finite element basis functions, we consider the Continuous Galerkin (CG) family (or the standard Lagrange family) with degree 1 or 2. See Table 4 for details.
Table 4: Configurations of data generation for different datasets
Dataset Configurations
Mesh $\phi_{i}^{\text{in}}$ $\phi_{i}^{\text{out}}$ $(N_{\text{train}},N_{\text{test}})$ $r$ $N_{\text{grad}}$ $s$
Darcy flow $64\times 64$, right CG (1D, deg=1) CG (1D, deg=1) $(1500,500)$ $16$ $16$ $10$
Hyperelasticity $64\times 64$, right CG (1D, deg=1) CG (2D, deg=1) $(1500,500)$ $16$ $16$ $10$
Navier–Stokes $64\times 64$, crossed CG (1D, deg=1) $u$: CG (2D, deg=2), $p$: CG (1D, deg=1) $(1500,500)$ $16$ $16$ $10$
We generate $N_{\text{train}}=1500$ and $N_{\text{test}}=500$ input-output pairs $(m^{(i)},u^{(i)})$ for training and testing, respectively. We compute first $r=16$ KLE basis and ASM basis using double pass randomized algorithm with an oversampling parameter $s$ of $10$. In the computation of ASM basis, we use $N_{\text{grad}}=16$ samples for the Monte Carlo estimate of the action of operator $\mathcal{H}$ in Equation 13. In our method, we formulate the different labels into arrays with the shape as follows
  • Evaluation labels: ($N_{\text{train}}$, $N_{x}$, $N_{u}$)
  • Derivative $m$ labels: ($N_{\text{train}}$, $N_{x}$, $r$, $N_{u}$)
  • Derivative $x$ labels: ($N_{\text{train}}$, $N_{x}$, $d$, $N_{u}$)
Recall that $N_{\text{train}}$ is the number of functions used for training, $N_{x}$ is the number of nodes of the mesh, $N_{u}$ is the dimension of output function, $r$ is the number of reduced basis, and $d$ is the dimension of the domain.
B.2 Training details
For the vanilla DeepONet, we parameterize the branch net using a CNN and the trunk net using a ResNet. For the reduced basis DeepONet, both the branch net and trunk net are parameterized using ResNets. We train each model for 1000 epochs. See Table 5, Table 6 and Table 7 for details about other hyperparameters we have used. When the loss functions comprises more than two type of losses, we enhance the model’s performance and training stability by incorporating the self-adaptive learning rate annealing algorithm [19], with an update frequency of $100$ and a moving average parameter of $0.9$. This algorithm automatically adjusts the loss weights $\{\lambda_{i}\}_{i=1}^{3}$ in Equation 5 at a minimal computational cost. Additionally, we standardize the inputs and labels before training.
Table 5: Training details for the vanilla DeepONet
Dataset
Darcy flow Hyperelasticity Navier–Stokes
Branch net CNN (7 layers, 128 out dim, ReLU) CNN (7 layers, 128 out dim, ReLU) CNN (7 layers, 128 out dim, ReLU)
Trunk net ResNet (7 layers, 128 out dim, ReLU) ResNet (7 layers, 256 out dim, ReLU) ResNet (7 layers, 256 out dim, ReLU)
Initialization Kaiming Uniform Kaiming Uniform Kaiming Uniform
(lr, weight decay) $(0.001,10^{-6})$ $(0.001,10^{-6})$ $(0.001,10^{-5})$
(gamma, step size) $(0.5,10)$ $(0.6,10)$ $(0.7,10)$
Table 6: Training details for the DE-DeepONet (KLE basis + dm labels + dx labels)
Dataset
Darcy Hyperelasticity Navier–Stokes
Branch net ResNet (4 layers, 128 out dim, ELU) ResNet (4 layers, 128 out dim, ELU) ResNet (7 layers, 128 out dim, ELU)
Trunk net ResNet (4 layers, 128 out dim, ELU) ResNet (4 layers, 256 out dim, ELU) ResNet (7 layers, 256 out dim, ELU)
Initialization Kaiming Normal Kaiming Normal Kaiming Normal
(lr, weight decay) $(0.005,10^{-13})$ $(0.001,10^{-12})$ $(0.001,0)$
(gamma, step size) $(0.9,10)$ $(0.99,50)$ $(0.9,20)$
$N_{x}^{\text{batch}}(=\alpha N_{x})$ $633\approx(0.15\times 65^{2})$ $422\approx(0.1\times 65^{2})$ $422\approx(0.1\times 65^{2})$
Table 7: Training details for the DE-DeepONet (ASM basis + dm labels + dx labels)
Dataset
Darcy Hyperelasticity Navier–Stokes
Branch net ResNet (4 layers, 128 out dim, ELU) ResNet (4 layers, 128 out dim, ELU) ResNet (7 layers, 128 out dim, ELU)
Trunk net ResNet (4 layers, 128 out dim, ELU) ResNet (4 layers, 256 out dim, ELU) ResNet (7 layers, 256 out dim, ELU)
Initialization Kaiming Normal Kaiming Normal Kaiming Normal
(lr, weight decay) $(0.005,10^{-13})$ $(0.001,10^{-12})$ $(0.001,10^{-12})$
(gamma, step size) $(0.9,10)$ $(0.99,50)$ $(0.9,10)$
$N_{x}^{\text{batch}}(=\alpha N_{x})$ $633\approx(0.15\times 65^{2})$ $422\approx(0.1\times 65^{2})$ $422\approx(0.1\times 65^{2})$
B.3 Test errors
We present the comparisons of relative error on test dataset (using the metrics in Section 4.3) of different methods considered in this work in Tables 8, LABEL:, 9, LABEL: and 10 for the Darcy flow equation, Tables 11, LABEL:, 12, LABEL: and 13 for the hyperelasticity equation, Tables 14, LABEL:, 15, LABEL: and 16 for the Navier–Stokes equations. For visualization, we provide the mean relative error plots in Figure 8 for the Darcy flow equation, Figure 9 for the hyperelasticity equation, and Figures 10, LABEL: and 4 for the Navier–Stokes equations.
Table 8: Darcy $5$ trails test relative $L^{2}(\Omega)$ error: mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 48.19 (5.53) 23.86 (0.77) 15.31 (0.75) 6.28 (0.33)
KLE 67.13 (4.94) 27.51 (0.95) 11.06 (0.15) 7.66 (0.09)
KLE + dm 24.40 (1.50) 10.46 (0.37) 4.89 (0.04) 4.44 (0.01)
KLE + dm + dx 22.34 (1.08) 9.39 (0.27) 4.87 (0.04) 4.47 (0.01)
ASM 63.78 (4.01) 26.26 (0.76) 10.76 (0.33) 6.87 (0.09)
ASM + dm 24.54 (1.26) 8.95 (0.18) 4.54 (0.03) 4.06 (0.02)
ASM + dm + dx 22.12 (1.03) 8.08 (0.24) 4.57 (0.02) 4.12 (0.01)
Table 9: Darcy $5$ trails test relative $H^{1}(\Omega)$ error: mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 83.11 (3.89) 50.02 (1.93) 35.34 (1.12) 20.69 (0.36)
KLE 85.88 (7.92) 44.78 (1.22) 24.75 (0.23) 21.60 (0.12)
KLE + dm 48.27 (1.65) 27.16 (0.74) 17.44 (0.15) 15.98 (0.07)
KLE + dm + dx 39.11 (1.67) 22.57 (0.31) 16.38 (0.08) 15.41 (0.03)
ASM 82.75 (5.79) 42.62 (1.10) 24.22 (0.25) 20.43 (0.19)
ASM + dm 41.07 (1.30) 24.59 (0.33) 17.03 (0.09) 15.67 (0.08)
ASM + dm + dx 35.62 (0.45) 21.27 (0.25) 16.06 (0.07) 15.06 (0.04)
Table 10: Darcy $5$ trails test relative $Fro$ error (for dm): mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 86.27 (3.25) 61.71 (2.02) 45.89 (2.89) 25.89 (2.32)
KLE 119.02 (5.27) 65.16 (2.02) 29.90 (0.38) 23.18 (0.27)
KLE + dm 57.11 (1.82) 28.97 (1.09) 11.06 (0.20) 8.89 (0.03)
KLE + dm + dx 55.32 (1.81) 26.42 (0.85) 11.21 (0.19) 9.20 (0.04)
ASM 120.13 (6.00) 62.25 (2.26) 28.83 (0.59) 20.44 (0.14)
ASM + dm 61.02 (2.91) 24.45 (0.62) 10.68 (0.06) 8.82 (0.05)
ASM + dm + dx 56.74 (1.05) 22.59 (0.90) 10.86 (0.13) 9.20 (0.05)
Refer to caption Refer to caption Refer to caption
Figure 8: Mean relative errors over 5 trials with a varying number of training samples for the Darcy equation using different methods. Relative errors in the $L^{2}(\Omega)$ norm (left) and $H^{1}({\Omega})$ norm (middle) for the prediction of $u$. Right: Relative error in the Frobenius (Fro) norm for the prediction of $du(m;\omega_{1})$.
Table 11: Hyperelasticity $5$ trails test relative $L^{2}(\Omega)$ error: mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 45.89 (1.15) 33.28 (2.10) 17.64 (0.90) 7.82 (0.49)
KLE 80.54 (10.88) 29.29 (0.84) 12.81 (0.39) 9.14 (0.49)
KLE + dm 17.18 (1.23) 9.14 (0.76) 5.00 (0.13) 4.73 (0.34)
KLE + dm + dx 17.11 (1.50) 8.61 (0.60) 5.07 (0.21) 4.57 (0.26)
ASM 80.36 (6.16) 29.74 (3.25) 13.11 (1.15) 7.06 (0.23)
ASM + dm 21.81 (1.53) 8.71 (0.82) 4.61 (0.31) 3.47 (0.23)
ASM + dm + dx 20.24 (1.38) 8.58 (0.81) 4.90 (0.81) 3.61 (0.21)
Table 12: Hyperelasticity $5$ trails test relative $H^{1}(\Omega)$ error: mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 57.80 (1.50) 41.87 (1.85) 26.23 (0.85) 16.77 (0.26)
KLE 80.91 (7.29) 38.88 (0.71) 21.73 (0.26) 17.46 (0.22)
KLE + dm 27.63 (0.91) 18.68 (0.17) 14.80 (0.13) 12.72 (0.27)
KLE + dm + dx 24.80 (1.01) 16.66 (0.49) 13.38 (0.31) 12.06 (0.26)
ASM 81.64 (5.59) 39.09 (1.88) 21.81 (0.50) 16.51 (0.15)
ASM + dm 31.15 (0.64) 18.76 (0.55) 14.56 (0.36) 13.02 (0.13)
ASM + dm + dx 27.38 (0.63) 16.64 (0.52) 13.26 (0.21) 11.92 (0.25)
Table 13: Hyperelasticity $5$ trails test relative $Fro$ error (for dm): mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 102.82 (0.86) 85.17 (5.59) 57.52 (3.76) 32.34 (2.33)
KLE 142.92 (11.76) 81.00 (3.57) 39.59 (0.74) 28.12 (0.83)
KLE + dm 50.93 (2.13) 25.82 (0.54) 15.39 (0.31) 13.08 (0.16)
KLE + dm + dx 43.09 (2.76) 22.16 (0.26) 12.93 (0.24) 10.91 (0.09)
ASM 138.26 (12.76) 81.01 (3.22) 37.78 (1.16) 22.47 (0.24)
ASM + dm 59.06 (1.67) 25.36 (1.41) 12.38 (0.29) 9.58 (0.05)
ASM + dm + dx 50.22 (2.48) 21.37 (0.84) 10.59 (0.12) 8.37 (0.20)
Refer to caption Refer to caption Refer to caption
Figure 9: Mean relative errors over 5 trials with a varying number of training samples for the hyperelasticity equation using different methods. Relative errors in the $L^{2}(\Omega)$ norm (left) and $H^{1}({\Omega})$ norm (middle) for the prediction of $u=(u_{1},u_{2})$. Right: Relative error in the Frobenius (Fro) norm for the prediction of $du(m;\omega_{1})=(du_{1}(m;\omega_{1}),du_{2}(m;\omega_{2}))$.
Table 14: Navier–Stokes $5$ trails test relative $L^{2}(\Omega)$ error: mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 41.28 (1.17) 26.64 (1.34) 16.18 (0.86) 7.90 (0.86)
KLE 64.44 (3.01) 28.19 (1.27) 16.39 (0.09) 14.87 (0.12)
KLE + dm 24.04 (1.13) 13.84 (0.25) 11.06 (0.09) 10.44 (0.04)
KLE + dm + dx 21.89 (0.80) 13.52 (0.29) 11.06 (0.05) 10.45 (0.03)
ASM 41.61 (1.12) 18.31 (1.03) 11.16 (0.54) 9.27 (0.21)
ASM + dm 21.33 (0.62) 10.16 (0.24) 6.81 (0.05) 6.25 (0.01)
ASM + dm + dx 21.88 (0.66) 10.69 (0.25) 7.09 (0.04) 6.27 (0.01)
Table 15: Navier–Stokes $5$ trails test relative $H^{1}(\Omega)$ error: mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 66.18 (1.42) 47.20 (2.02) 23.10 (0.80) 13.17 (0.58)
KLE 61.97 (1.44) 34.53 (1.12) 21.86 (0.19) 21.45 (0.10)
KLE + dm 51.81 (2.07) 31.57 (1.11) 18.29 (0.27) 15.40 (0.02)
KLE + dm + dx 35.54 (0.91) 27.28 (0.23) 18.53 (0.12) 15.42 (0.03)
ASM 51.19 (2.65) 28.48 (2.90) 16.48 (1.70) 14.10 (0.29)
ASM + dm 48.08 (1.27) 26.71 (0.54) 13.28 (0.37) 10.57 (0.01)
ASM + dm + dx 47.78 (2.11) 27.28 (0.41) 17.87 (0.48) 10.91 (0.17)
Table 16: Navier–Stokes $5$ trails test relative $Fro$ error (for dm): mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 576.93 (42.09) 375.72 (51.98) 257.51 (25.55) 212.01 (42.70)
KLE 607.06 (18.07) 293.72 (15.57) 144.20 (4.31) 165.90 (4.86)
KLE + dm 199.44 (7.48) 113.26 (7.47) 68.60 (2.01) 56.65 (0.62)
KLE + dm + dx 197.82 (10.62) 113.70 (5.85) 69.96 (2.14) 58.49 (0.39)
ASM 195.21 (9.51) 78.99 (2.44) 50.55 (0.69) 49.05 (0.60)
ASM + dm 104.05 (1.11) 52.88 (1.51) 30.67 (0.16) 26.36 (0.08)
ASM + dm + dx 102.07 (1.80) 54.23 (1.54) 31.47 (0.18) 26.03 (0.07)
Refer to caption Refer to caption Refer to caption
Figure 10: Mean relative errors over 5 trials with a varying number of training samples for the Navier–Stokes equations using different methods. Relative errors in the $L^{2}(\Omega)$ norm (left) and $H^{1}({\Omega})$ norm (middle) for the prediction of the velocity-x. Right: Relative error in the Frobenius (Fro) norm for the prediction of the derivative of the velocity-x with respect to the parameter in the direction of the 128 samples drawn from the same distribution as parameter.
B.4 Reduced basis
We plot the eigenvalues of KLE basis and ASM basis in Figure 11. We present visualizations of the first three KLE and ASM bases in Figure 12.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 11: First 100 eigenvalues of KLE and ASM basis. Top row: Darcy; Middle row: Hyperelasticity; Bottom row: Navier–Stokes
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 12: First row: Darcy first $3$ KLE basis (similar patterns observed for Hyperelasticity and Navier–Stokes); Second row: Darcy first $3$ ASM basis; Third row: Hyperelasticity first $3$ ASM basis; Fourth row: Navier–Stokes first $3$ ASM basis.
To measure the error induced by the projection, we define the output reconstruction error as follows
$\frac{1}{N}\sum_{i=1}^{N}\frac{\|u(P_{r}m^{(i)})-u(m^{(i)})\|_{L^{2}}}{\|u(m^{(i)})\|_{L^{2}}},$
where $P_{r}$ is the rank $r$ linear projector. We provide the plots of the output reconstruction error vs number of reduced basis $r$ using KLE basis and ASM basis in Figure 13. We can see that using the ASM basis results in a lower output reconstruction error than the KLE basis across all three equations.
Refer to caption Refer to caption Refer to caption
Figure 13: Output reconstruction error using KLE and ASM basis. Left: Darcy; Middle: Hyperelasticity; Right: Navier–Stokes
B.5 Visualization of the ground truth and prediction
We present the comparisons of the ground truth of solution $u$, model prediction $\hat{u}$ using different methods, and absolute value of their difference $|u(x)-\hat{u}(x)|$ in Figure 14 for the Darcy flow equation,  Figures 16, LABEL: and 17 for the hyperelasticity equation, and Figures 20, LABEL: and 5 for the Navier–Stokes equations.
We present the comparisons of the ground truth of the derivative of $u$ with respect to $m$ in the direction of $\omega_{1}$, denoted as $du(m;\omega_{1})$, model prediction $d\hat{u}(m;\psi_{1}^{\text{in}})$ using different methods, and absolute value of their difference $|du(m;\omega_{1})-d\hat{u}(m;\omega_{1})|$ in Figure 15 for the Darcy flow equation,  Figures 18, LABEL: and 19 for the hyperelasticity equation, and Figures 21, LABEL: and 6 for the Navier–Stokes equations.
Figure 14: Darcy flow. Comparison of the predictions of $u$ with (16, 64, 256, 1024) training samples using different methods. First row: Vanilla DeepONet; Second row: KLE; Third row: ASM; Fourth row: KLE + dm +dx; Fifth row: ASM + dm + dx.
Figure 15: Darcy flow. Comparison of the predictions of directional derivative $du(m;\omega_{1})$ with (16, 64, 256, 1024) training samples using different methods. First row: Vanilla DeepONet; Second row: KLE; Third row: ASM; Fourth row: KLE + dm +dx; Fifth row: ASM + dm + dx.
Figure 16: Hyperelasticity. Comparison of the predictions of $u_{1}$ with (16, 64, 256, 1024) training samples using different methods. First row: Vanilla DeepONet; Second row: KLE; Third row: ASM; Fourth row: KLE + dm +dx; Fifth row: ASM + dm + dx.
Figure 17: Hyperelasticity. Comparison of the predictions of $u_{2}$ with (16, 64, 256, 1024) training samples using different methods. First row: Vanilla DeepONet; Second row: KLE; Third row: ASM; Fourth row: KLE + dm +dx; Fifth row: ASM + dm + dx.
Figure 18: Hyperelasticity. Comparison of the predictions of directional derivative $du_{1}(m;\omega_{1})$ with (16, 64, 256, 1024) training samples using different methods. First row: Vanilla DeepONet; Second row: KLE; Third row: ASM; Fourth row: KLE + dm +dx; Fifth row: ASM + dm + dx.
Figure 19: Hyperelasticity. Comparison of the predictions of directional derivative $du_{2}(m;\omega_{1})$ with (16, 64, 256, 1024) training samples using different methods. First row: Vanilla DeepONet; Second row: KLE; Third row: ASM; Fourth row: KLE + dm +dx; Fifth row: ASM + dm + dx.
Figure 20: Navier–Stokes. Comparison of the predictions of velocity-x with (16, 64, 256, 1024) training samples using different methods. First row: Vanilla DeepONet; Second row: KLE; Third row: ASM; Fourth row: KLE + dm +dx; Fifth row: ASM + dm + dx.
Figure 21: Navier–Stokes. Comparison of the predictions of directional derivative $du_{1}(m;\omega_{1})$ with (16, 64, 256, 1024) training samples using different methods. First row: Vanilla DeepONet; Second row: KLE; Third row: ASM; Fourth row: KLE + dm +dx; Fifth row: ASM + dm + dx.